OpenStudy (anonymous):

Solve. x2 - 3x = 18. (3 points) Answer x = 2 and x = -2 x = 6 and x = 2 x = 3 and x = -6 x = 6 and x = -3

hartnn (hartnn):

your answer ?

hartnn (hartnn):

i would have tried substituting the options....

OpenStudy (anonymous):

It's not B for sure

hartnn (hartnn):

correct, not B

OpenStudy (anonymous):

C

hartnn (hartnn):

sure ? if u put x=3 3^2 - 3*3 =.... ?

OpenStudy (anonymous):

oh...

hartnn (hartnn):

so, not C

OpenStudy (anonymous):

81

hartnn (hartnn):

3^2 -3*3 = 9-9 =0 and not 18

OpenStudy (anonymous):

hmmm...

hartnn (hartnn):

try D

OpenStudy (anonymous):

Ok... :)

OpenStudy (anonymous):

It worked

hartnn (hartnn):

yes, its D :)

OpenStudy (anonymous):

YAYYYYYYYYYYYYY:))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))